# Emancipatory Information Retrieval

### v1

@@ Abstract
Our world today is facing a confluence of several mutually reinforcing crises each of which intersects with concerns of social justice and emancipation. This paper is a provocation for the role of computer-mediated information access in our emancipatory struggles. We define emancipatory information retrieval as the study and development of information access methods that challenge various forms of human oppression, and situates its activities within broader collective emancipatory praxis. The term “emancipatory” here signifies the moral concerns of universal humanization of all peoples and the elimination of oppression to create the conditions under which we can collectively flourish. To develop an emancipatory research agenda for IR, in this paper we speculate about the practices that the community can adopt, enumerate some of the projects that the field should undertake, and discuss provocations to spark new ideas and directions for research. We challenge the field of information retrieval (IR) research to embrace humanistic values and commit to universal emancipation and social justice as part of our research.
@@ 3.2 Emancipatory projects in IR
Safeguard against discrimination.
Emancipatory IR should concern itself with ensuring that IR systems do not lead to inequitable outcomes which may further reinscribe historical sociopolitical and economic oppression and exploitation. Over the last several years, there has been several works on fairness in IR. These have typically focused on fairness of quality-of-service (e.g.[89, 90, 97, 142]) and fairness of exposure (e.g.[20, 113, 114, 41, 145, 103, 141]). However, there is less clarity on how much this body of fairness research has practically influenced material changes in popular IR platforms [91]. Fairness research in the emancipatory context must move beyond just formalization and theorization of fairness and must contend with how this research opposes material inequities in the real world. This includes also contending with the politics of classification [38] and how certain practices in fairness research, such as in context of gender fairness [105], may themselves be harmful.
We must also hold institutions that own and operate IR platforms accountable to measurable equitable and non-discriminatory outcomes. If an institution invests in fairness research but do not operationalize them in their products, then that research only serves to launder institutional reputation and is counter to our real emancipatory goals. To hold institutions accountable, our research agenda should also include challenges of auditing IR platforms and also explore other regulatory, sociopolitical, and technological mechanisms of accountability that is grounded in the recognition of the structural power inequities that make it challenging to do so.

### v2

@@ Abstract
Our world today is facing a confluence of several mutually reinforcing crises each of which intersects with concerns of social justice and emancipation. This paper is a provocation for the role of computer-mediated information access in our emancipatory struggles. We define emancipatory information retrieval as the study and development of information access methods that challenge various forms of human oppression, and situates its activities within broader collective emancipatory praxis. The term "emancipatory" here signifies the moral concerns of universal humanization of all peoples and the elimination of oppression to create the conditions under which we can collectively flourish. To develop an emancipatory research agenda for IR, in this paper we speculate about the practices that the community can adopt, enumerate some of the projects that the field should undertake, and discuss provocations to spark new ideas and directions for research. We challenge the field of information retrieval (IR) research to embrace humanistic values and commit to universal emancipation and social justice as part of our research.
@@ 3.2 Emancipatory projects in IR
Safeguard against dehumanization and inequitable outcomes.
Emancipatory IR should concern itself with ensuring that IR systems do not lead to representational harms (e.g., denigration, stereotyping, and erasure) of historically marginalized peoples and inequitable outcomes which may further reinscribe historical sociopolitical and economic oppression and exploitation. Over the last several years, there has been several works on fairness in IR. These have typically focused on fairness of quality-of-service (e.g.[89, 90, 97, 142]) and fairness of exposure (e.g.[20, 113, 114, 41, 145, 103, 141]). However, there is less clarity on how much this body of fairness research has practically influenced material changes in popular IR platforms [91]. Fairness research in the emancipatory context must move beyond just formalization and theorization of fairness; it must recognize and be grounded in the understanding of historical injustices and structural violence, and must ensure that this research opposes dehumanization and material inequities in the real world. This includes also contending with the politics of classification [38] and critiquing practices in fairness research, such as in context of gender fairness [105], that may themselves perpetuate further harm.
We must also hold institutions that own and operate IR platforms accountable to measurable equitable and humanistic outcomes. If an institution invests in fairness research but do not operationalize them in their products, then that research only serves to launder institutional reputation and is counter to our real emancipatory goals. To hold institutions accountable, our research agenda should also include challenges of auditing IR platforms and also explore other regulatory, sociopolitical, and technological mechanisms of accountability that is grounded in the recognition of the structural power inequities that make it challenging to do so.